OpenStudy (anonymous):

solve the given equation. 6+8e^2x=27 ?

OpenStudy (anonymous):

@ganeshie8 ??

ganeshie8 (ganeshie8):

subtract 6 both sides

ganeshie8 (ganeshie8):

8e^2x = 21 divide 8 both sides e^2x = 21/8

ganeshie8 (ganeshie8):

fine so far ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so whats the answer @ganeshie8

ganeshie8 (ganeshie8):

wolfram gives u answers you dont need me friend :)

ganeshie8 (ganeshie8):

click on "Approximate form"